# Data-Driven Lipschitz Continuity: A Cost-Effective Approach to Improve Adversarial Robustness

### v1

@@ Abstract
The security and robustness of deep neural networks (DNNs) have become increasingly concerning. This paper aims to provide both a theoretical foundation and a practical solution to ensure the reliability of DNNs. We explore the concept of Lipschitz continuity to certify the robustness of DNNs against adversarial attacks, which aim to mislead the network with adding imperceptible perturbations into inputs. We propose a novel algorithm that remaps the input domain into a constrained range, reducing the Lipschitz constant and potentially enhancing robustness. Unlike existing adversarially trained models, where robustness is enhanced by introducing additional examples from other datasets or generative models, our method is almost cost-free as it can be integrated with existing models without requiring re-training. Experimental results demonstrate the generalizability of our method, as it can be combined with various models and achieve enhancements in robustness. Furthermore, our method achieves the best robust accuracy for CIFAR10, CIFAR100, and ImageNet datasets on the RobustBench leaderboard.
@@ 1 Introduction
Deep neural networks (DNNs) have demonstrated promising results across various tasks [1, 2], prompting concerns about AI security as these networks are increasingly deployed in our daily lives. A single erroneous prediction could lead to catastrophic consequences. For example, the Overload attack can significantly inflate the inference time of detecting objects for self-driving systems [3], while even minor typos in input prompts can cause large language models (LLMs) to produce unexpected responses [4].
The focus of this paper is to design robust DNNs that can defend against adversarial attacks, which aim to create perturbations in inputs that are imperceptible to humans but can mislead DNNs. Previous studies have revealed the existence of adversarial examples in diverse domains, such as image pixels [5], audio data [6], and textual content [7]. Consequently, exploring the vulnerabilities of DNNs and developing theoretically grounded explainable AI is crucial for ensuring the reliability of DNN-based applications.
@@
Adversarial training [8] has proven to be an effective strategy for enhancing the robustness of DNNs. It achieves this by generating adversarial examples on the fly during the training phase and optimizing the model’s weights to minimize the losses caused by these examples. Recent studies have shown that robustness can be further improved by introducing additional examples from other datasets [9] or using generative models [10, 11] to cover low-frequency data. Despite the promising improvements in robustness, training costs increase significantly due to the demand for additional data, which can be up to 20 to 100 times larger than the original dataset. This poses a trade-off between training cost and robustness. The concern over high computational costs becomes a significant obstacle in deploying robust DNN-based applications, especially in fields like medicine, autonomous driving systems, and other areas where human lives are at stake.
In this paper, we explore how robustness is certified by the theorem of Lipschitz continuity, which theoretically gauges how much outputs are amplified by the perturbations. However, we argue that the set of observed data is finite and cannot cover the entire real data space, leading to an overestimation of the Lipschitz constant derived from the theorem. Therefore, we propose an algorithm that can remap the input domain into a constrained range, resulting in a Lipschitz constant for the modified function that is less than or equal to the Lipschitz constant of the original function, thus potentially enhancing robustness. Our key contributions are outlined as follows:
  • We introduce the concept of the empirical Lipschitz constant, which can more precisely reflect the robustness of the corresponding observed data. Compared with the original definition of Lipschitz constant, the empirical value is derived from a set of observed data, thereby eliminating the influence of space that is never drawn from real data. As illustrated in Figure 1, we prove that any function that can be formulated as a linear system, when combined with our proposed function to remap the input domain of a specific layer to a constrained range, can reduce its empirical Lipschitz constant, resulting in better robustness.
  • The proposed function can enhance the robustness of adversarially trained models with minimal additional costs. Specifically, it introduces only one parameter, the value of which can be determined by scanning an observed data once without the need for re-training or fine-tuning.
  • The experimental results suggest that our method can be combined with various existing methods and gain robustness improvements. Besides, our method achieves the best robust accuracy against adversarial examples generated by AutoAttack [12], a state-of-the-art ensemble attack, for CIFAR10, CIFAR100, and ImageNet datasets on the RobustBench leaderboard [13]. By assessing accuracy against adaptive attacks, transfer attacks, and evaluation methods for validating obfuscated gradients [14, 15], we believe that the proposed algorithm should not cause robustness to be overestimated.
Figure 1: The empirical Lipschitz constant of specific layers that can be represented by linear systems, such as convolutional or fully connected layers, can be reduced by remapping their input domain to a constrained range.
@@
The rest of this paper is organized as follows. Section 2 introduces the background on adversarial attacks and adversarial training. Section 3 presents the theoretical proof of how robustness is enhanced by manipulating the domain of linear functions and introduces the proposed algorithm. Section 4 shows the experimental results, including the comparisons among related works, and ablation studies on various hyper-parameters, combination with different activation functions and gradient masking verification. The last section is our conclusion.
@@ 2.1 Adversarial Attacks
Adversarial attacks aim to inject tiny perturbations into inputs, causing victim DNNs to output incorrect predictions with high confidence [16]. These attacks have been observed in numerous vision applications [17, 18, 19, 20]. Furthermore, these tiny perturbations can be embedded not only in image pixels but also in textual contexts [21, 22], audio space [23], and other fields [24]. Some research has shown how adversarial attacks threaten real applications [25, 26, 27, 28]. Investigating the vulnerability of DNNs and theoretically avoiding adversarial examples when optimizing model weights or designing architecture is an ongoing challenge.
Depending on the amount of information the attacker has access to, adversarial attacks can be divided into two types: white-box attacks and black-box attacks. For white-box attacks, all information about the victim models is public. Attackers can craft adversarial examples through the gradient direction, which is usually opposite to the direction in which the model weights were optimized during the training phase [8, 12]. Although this scenario is unrealistic, this type of research could lead to the design of more reliable models in the future. Conversely, for black-box attacks, the only information leaked to the attacker is the output prediction of the victim model. Adversarial examples can be generated by random search [29], discrepancies in outputs [30], or transferability from models with similar architectures [31, 32]. The purpose of black-box attacks is to study the risk of the victim models being attacked in real-world application scenarios.
@@ 2.2 Defensive Strategies
Adversarial training is a defensive strategy that aims to find optimal weights against adversarial attacks [8, 33, 34]. It achieves this by generating adversarial examples on the fly during the training phase and optimizing the model’s weights to minimize the losses caused by these examples. Despite the superior robustness achieved by adversarial training, the associated training costs of adversarially trained models are generally ten times more expensive than those of models trained utilizing a standard policy. The concern over high computational costs becomes a significant obstacle in deploying DNN-based applications.
@@ 3 Methodology
3.1 Motivation
In this paper, we approach robustness from a theoretical perspective, aiming to demonstrate that all risks posed by adversarial examples are limited while minimizing the additional costs associated with improving robustness. Our evaluation is conducted under the white-box scenario, where the target model is capable of defending against various types of known adversarial attacks, including white-box attacks [8, 18, 12], black-box attacks [43], and transfer attacks [44, 45]. Additionally, we conduct a set of experiments to verify that gradient masking [14] does not occur in our method and to ensure that robustness is not overestimated.
@@ 3.2 Lipschitz Continuity
$\frac{||f(x^{\text{adv}})-f(x)||_{2}}{||x^{\text{adv}}-x||_{2}}=\frac{(||Wx^{\text{adv}}+b)-(Wx+b)||_{2}}{||\delta||_{2}}=\frac{||W\delta||_{2}}{||\delta||_{2}},$ (4)
@@
$\min_{W}\max_{\delta\neq 0,\delta\in\mathbb{R}^{m}}\frac{||W\delta||_{2}}{||\delta||_{2}}=\min_{W}\sigma_{\text{max}}(W),$ (5)
@@
Rather that minimizing the objective directly, Gershgorin circle theorem provides an alternative solution to estimate the robustness of the given linear system.
@@
3.3 Forged Function
(a) ConvNet
(b) Transformer
We argue that the largest singular value provides a loose bound for the Lipschitz constant. To precisely reflect the robustness of the corresponding observed data, we define the empirical Lipschitz constant that eliminates the influence of space that is never drawn from real data.
Definition 1.
Empirical Lipschitz constant:
$\max_{\delta\neq 0,x\in\mathcal{S}}\frac{||Wx||_{2}}{||x||_{2}}\quad\forall x\in\mathcal{S},$ (8)
where $\mathcal{S}$ is an observed dataset. As can be seen, the empirical Lipschitz constant on the finite dataset is less than or equal to its Lipschitz constant derived from the theorem.
Based on Definition 1, we can build robust models by manipulating the output ranges of individual layers, thereby restricting the input domain of the next layer. If input vectors do not align with the direction of the eigenvector with the largest eigenvalue, the empirical constant should be bounded. Therefore, we proposed a forged function defined as follows:
$f^{\text{forge}}(x)=\begin{cases}0\quad&\text{if}\quad|x|\leq c^{\text{th}}_{i},\\ x\quad&\text{otherwise},\end{cases}$ (9)
where $c^{\text{th}}_{i}$ is a threshold for the $i$-th layer. Compared with the original functions, the range of the forged function is suppressed if its value is less than the threshold. When $c^{\text{th}}_{i}$ is set to 00, the forged function degrades into the original function.
The forged function aim to reduce the empirical Lipschitz constant of the layers that can be represented as linear systems by remapping the input domain of these layers into a constrained set. Figure 2 provides a visual representation of potential insertion points for the forged function, while maintaining the integrity of other layers. For the ResNet architecture, the forged function is placed before the convolutional layers in each residual block. Similarly, for vision transformer architectures, the structure of MLP layers is adapted to seamlessly integrate the forged function.
Here is the proof that the largest eigenvalue can be shrunk by the forged function. Let $W$ be the weight of the target layer, which can be represented by an $m\times n$ matrix, and t be the input vector. Without loss of generality, we assume that $A=W^{\dagger}W$ and $f^{\text{forge}}(\textbf{t})$ is defined as:
$f^{\text{forge}}(t_{i})=\begin{cases}0\quad i\leq k\\ t_{i}\quad\text{otherwise},\end{cases}$ (10)
where $t_{i}$ is the $i$-th element of t and $k$ is a positive number.
@@ 3.3 Forged Function
Since the first $k$ entries of the vector t are replaced with zeros, above condition can be achieved by replacing the corresponding column vectors of the matrix $A$ with zero vectors. Therefore, the entries of $A^{\prime}$ are formulated as
$a^{\prime}_{ij}=\begin{cases}0\quad j\leq k\\ a_{ij}\quad\text{otherwise}.\end{cases}$ (12)
The matrix $A$ is a positive semidefinite matrix, implying that the diagonal entries are non-negative. Moreover, with the entry representation of $A^{\prime}$ in (12), we observe that modifications are only applied to the first $k$ columns, while the rest remain unchanged. Combining the Gershgorin Circle Theorem, we know that the centers of the first $k$ disks of the matrix $A^{\prime}$ are shifted towards zero. Additionally, the radii of all disks, the absolute values of the off-diagonal entries in $A^{\prime}$, are shrunk. Consequently, the upper bound of the largest eigenvalue of the matrix $A^{\prime}$ is tighter compared to that of the original matrix $A$. ∎
Notably, the outputs of $f^{\text{forge}}(\textbf{t})$ vary depending on the inputs, resulting in each input having its own $A^{\prime}$. The upper bound of the largest eigenvalue of each matrix $A^{\prime}$ must be not greater than the largest eigenvalue of the matrix $A$. With Lamma 2, a precise upper bound of the largest eigenvalue can be obtained by feeding a set of observed images. On the contrary, there might be cases in which solving the minimization problem in (5) leads to the theoretical largest eigenvalue being minimized, but the empirical Lipschitz constant remains unchanged.
The choice of a proper $c^{\text{th}}_{i}$ is a crucial factor in reducing the largest eigenvalue. In this paper, we propose obtaining the value of $c^{\text{th}}_{i}$ through the following equation:
@@
where $\mathcal{S}$ can include all or a subset of images in the training set, $c^{r}$ is a positive number and $F_{1\rightarrow i}(x)$ represents the output of the $i$-th layer. Specifically, each layer has its own $c^{\text{th}}_{i}$, but they share the same hyper-parameter $c^{r}$. Algorithm 1 specifies the implementation details of the forged function. The variable $b$ is used to store the maximum value that appeared in $\mathcal{S}$, as defined in (13), and is initialized during construction. Similar to the implementation of the batchnorm layer, the behavior is depended on the mode configuration. When the mode is set to tracking mode, the variable $b$ is updated accordingly, and the input is set to the output without any modification. Conversely, when the mode is set to inference mode, the value of $b$ is frozen, but the input is updated as defined by (9). By default, the mode is set to inference, and the values of $b$ and $c^{r}$ are zero, respectively. As a result, the set $\mathcal{M}$ is empty, and the algorithm is degraded to the identical function.
It is worth emphasizing that by feeding all images in the set $\mathcal{S}$ once in track mode beforehand, the value of $c^{\text{th}}_{i}$ can be obtained. The elements satisfying the constraints are appropriately deactivated during inference. Notably, this operation does not necessitate gradient computations and incurs minimal time consumption, typically only a few minutes, even when executed on commonly used GPUs. In comparison to adversarial training, this process is nearly cost-free. The overall procedure shares many similarities with post-pruning techniques. Nevertheless, we posit that the proposed function is very similar to the ReLU function, as it suppresses the output values within a specific range, but the defined range in the forged function is adaptive to the observed dataset.
@@
1:  require: Input x, Mode $m$
@@
5:     =$\mathcal{M}=\left\{x\middle|\text{abs}(x)\leq c^{r}b\right\}$
6:     for all $s\in\mathcal{M}$ do
7:        $s=0$
8:     end for
9:  end if
10:  return x

### v2

@@ Abstract
As deep neural networks (DNNs) are increasingly deployed in sensitive applications, ensuring their security and robustness has become critical. A major threat to DNNs arises from adversarial attacks, where small input perturbations can lead to incorrect predictions. Recent advances in adversarial training improve robustness by incorporating additional examples from external datasets or generative models. However, these methods often incur high computational costs, limiting their practicality and hindering real-world deployment. In this paper, we propose a cost-efficient alternative based on Lipschitz continuity that achieves robustness comparable to models trained with extensive supplementary data. Unlike conventional adversarial training, our method requires only a single pass over the dataset without gradient estimation, making it highly efficient. Furthermore, our method can integrate seamlessly with existing adversarial training frameworks and enhances the robustness of models without requiring extra generative data. Experimental results show that our approach not only reduces computational overhead but also maintains or improves the defensive capabilities of robust neural networks. This work opens a promising direction for developing practical, scalable defenses against adversarial attacks.
@@ 1 Introduction
Deep neural networks (DNNs) have achieved remarkable success across a wide range of tasks [29, 38], driving their adoption in increasingly sensitive applications. As a result, concerns over the security and robustness of these models have become more urgent, as even a single erroneous prediction can have serious consequences. For example, the Overload attack can significantly inflate the inference time of detecting objects for edge devices [12], and even minor typos in prompts can cause large language models to behave unpredictably [55]. Among the most critical threats are adversarial attacks, which add imperceptible perturbations to inputs, leading DNNs to make incorrect predictions. These vulnerabilities have been demonstrated across various domains [40, 6, 31], highlighting the urgent need for more robust and explainable AI models.
@@
Adversarial training [34] has proven to be an effective strategy for enhancing the robustness of DNNs by generating adversarial examples during training and optimizing model weights to minimize the losses caused by these examples. Despite promising results, no certified models have been deployed commercially due to prohibitive computational requirements. Recent studies have shown that robustness can be further improved by introducing additional examples from other datasets [8] or using generative models [24, 43] to cover low-frequency data. However, these approaches require datasets 100 or more times larger than the original, which significantly escalates computational demands by orders of magnitude. This creates a critical trade-off between training cost and robustness, presenting a significant obstacle to deploying robust DNN-based applications commercially are at stake.
The focus of this paper is to explore an alternative approach that reduces training cost while achieving robustness comparable to state-of-the-art adversarial training frameworks. To address this issue, we first revisit the theorem of Lipschitz continuity, which gauges how much outputs are amplified by perturbations. Unlike certified training [35] or Lipschitz-constrained methods [56] approaches that primarily aim to tighten upper bounds but often face scalability issues with larger models and datasets, our approach takes a fundamentally different direction. We prove that remapping the input domain of any layer that can be represented as a linear system to a constrained range results in a bounded Lipschitz constant and robustness improvements when specific conditions are met.
The key insight is that the forged function that reshapes the input domain of a linear system should maintain high similarity between the output of the transformed layers and the original, according to specific metrics. We propose a data-driven algorithm that constructs the optimal function with only a single scan over the dataset and no need for gradient estimation, making it highly efficient. This approach stands in contrast to prior methods that typically reshape weight distributions through regularization or by introducing complex objectives, which increases the computational cost of seeking suitable hyper-parameters.
Our key contributions are outlined as follows:
  • Cost-Effective Robustness Enhancement: We propose a method that enhance the robustness of existing adversarially trained methods with minimal additional costs. Compared to previous works, where improved robustness relied on introducing more generative data during the adversarial training phase, our method can be combined with existing models without re-training or fine-tuning. From a different perspective, for models with similar robustness, those integrated with our approach require significantly lower training cost.
  • Theoretical Foundation: We provide a theoretical proof that, under specific conditions, remapping the input domain of any linear system to a constrained range leads to a bounded Lipschitz constant and improved robustness. This theoretical foundation distinguishes our work from previous Lipschitz-based approaches that focus primarily on constraining the network during training.
  • Inference-Time Optimization: The proposed method is well-suited for inference-time model optimization and is almost cost-free. The function introduces only one parameter, the value of which can be determined by scanning observed data once without requiring gradient computations or fine-tuning. This makes our approach particularly valuable for practical applications where computational resources are limited.
  • State-of-the-Art Results: Experimental results demonstrate that our method can be combined with various existing methods to achieve meaningful robustness improvements. Notably, models such as LTD [10] and DefEAT [11], when combined with our method, achieve performance on par with RST-AWP [45], a method that relies on extra data during training, against against adversarial examples generated by AutoAttack [17], a state-of-the-art ensemble attack, for CIFAR10, CIFAR100, and ImageNet datasets on the RobustBench leaderboard [18]. This finding supports the hypothesis that certain data samples may be safely omitted without sacrificing robustness, reinforcing the potential of developing data selection algorithms for adversarial training to reduce training cost without compromising effectiveness.
@@
The rest of this paper is organized as follows. Section 2 introduces the background on adversarial attacks and adversarial training. Section 3 presents the theoretical proof of how robustness is enhanced by manipulating the domain of linear functions and introduces the proposed algorithm. Section 4 shows the experimental results, including the comparisons among related works. Due to page limitations, ablation studies on various hyper-parameters, combination with different activation functions and gradient masking verification are left in Appendix. The last section is our conclusion.
@@ 2.1 Adversarial Attacks
Adversarial attacks aim to inject tiny perturbations into inputs, causing victim DNNs to output incorrect predictions with high confidence [14]. These attacks have been observed in numerous vision applications [22, 15, 42, 49]. Furthermore, these tiny perturbations can be embedded not only in image pixels but also in textual contexts [30, 48], audio space [46], and other fields [27]. Some research has shown how adversarial attacks threaten real applications [47, 28, 20, 44]. Investigating the vulnerability of DNNs theoretically and designing robust DNNs is an ongoing challenge.
Adversarial attacks can be classified into two types based on the amount of information the attacker has access to: white-box attacks and black-box attacks. In the case of white-box attacks, the attacker has full access to all information about the victim model. Methods such as PGD attack [34] and AutoAttack [17] generate adversarial examples by leveraging the gradient direction. Although this scenario is often unrealistic in practical settings, research in this area is valuable for developing more robust models in the future.
In contrast, black-box attacks assess the risks of adversarial attacks in real-world scenarios where attackers have limited access to the model’s internals. Square Attack [2] and ZO-NGD [54] demonstrate that adversarial examples can be generated by querying the model’s output predictions. Additionally, adversarial examples can be crafted through transferability, where models with similar architectures are used to generate adversarial examples that remain effective across different models [41, 13].
@@ 2.2 Defensive Strategies
Adversarial training is a defensive strategy that aims to find optimal weights against adversarial attacks. It achieves this by generating adversarial examples on the fly during the training phase and optimizing the model’s weights to minimize the losses caused by these examples. Despite the superior robustness achieved by adversarial training, the associated training costs of adversarially trained models are generally ten times more expensive than those of models trained utilizing a standard policy. The concern over high computational costs becomes a significant obstacle in deploying DNN-based applications.
@@ 3 Methodology
The main goal of this paper is to enhance the robustness of adversarially trained models by leveraging Lipschitz continuity theory to remap the input domain. We propose a data-driven approach that aachieves state-of-the-art performance with minimal extra computational cost, effectively balancing model resilience and training efficiency.
@@ 3.1 Lipschitz Continuity
[…]
@@
[…]
@@
3.2 Forged Function
One approach to improving the robustness of a neural network is to minimize the objective function defined in (5) to obtain an optimal weight matrix. However, we argue that the largest singular value provides only a loose bound for the Lipschitz constant, as adversarial examples typically lie within an $\epsilon$-ball with a small radius, rather than spanning the entire input space. This observation suggests an alternative strategy: remapping the input domain to a constrained range, which effectively shrinks the reachable space. The restricted set should suppress values in specific dimensions most affected by adversarial examples, thereby minimizing the distribution mismatch between natural and adversarial data, all while maintaining accuracy. Therefore, the problem can be decomposed into two parts: (a) identifying an appropriate constrained range, and (b) transforming it into a corresponding weight matrix that ensures a bounded Lipschitz constant.
We propos a forged function for each layer defined as follows:
$f^{\text{forge}}_{i}(x)=\begin{cases}xf^{\text{masked}}_{i}(x)\quad&\text{if}\quad|x|\leq c^{\text{th}}_{i},\\ x\quad&\text{otherwise},\end{cases}$ (7)
where $c^{\text{th}}_{i}$ is a threshold for the $i$-th layer; $f^{\text{masked}}_{i}(x)$ is a scaling function with values in the range $[0,1]$. Compared with the original input, the range of the forged function is suppressed if its value is less than the threshold. When $c^{\text{th}}_{i}$ is set to 0, the forged function degrades into the original function.
The simplest definition of scaling function is a step function:
$f^{\text{step}}(x)=\begin{cases}0\quad&|x|\leq c^{\text{th}}_{i},\\ 1\quad&\text{otherwise}.\end{cases}$ (8)
However, discontinuous functions might cause unreliable gradients, potentially exposing the defense to BPDA attacks [3]. Therefore, we propose a smooth function that gradually reduces the value to 0, defined as:
$f^{\text{smooth}}=\frac{1}{1+\exp(-ax+b)},$ (9)
where $a$ controls the slope, and $b$ determines the location of the transition zone. Alternatively, a piecewise linear function can simplify this behavior:
$f^{\text{piecewise}}(x)=\begin{cases}0\quad&|x|\leq dc^{\text{th}}_{i},\\ \frac{|x|-dc^{\text{th}}_{i}}{(1-d)c^{\text{th}}_{i}}\quad&dc^{\text{th}}_{i}\leq|x|\leq c^{\text{th}}_{i},\\ 1\quad&\text{otherwise},\end{cases}$ (10)
where $d$ is within the range $[0,1]$. In (10), the transition zone is represented by a linear function, and as a result, the composed $f^{\text{forge}}$ in (7) becomes a quadratic function within this zone.
(a) ConvNet
(b) Transformer
Figure 1 provides a visual representation of potential insertion points for the forged function, while maintaining the integrity of other layers. For the ResNet architecture, the forged function is placed before the convolutional layers in each residual block. Similarly, for vision transformer architectures, the structure of MLP layers is adapted to seamlessly integrate the forged function. However, although the Query (Q), Key (K), and Value (V) computations in the attention layers involve linear projections, we do not apply the forged function to these components due to the more complex implementation involved.
To understand how the forged function influences the Lipschitz constant, we begin by recalling the Gershgorin Circle Theorem:
@@ 3.2 Forged Function
Now, let $W$ be the weight of the target layer, which can be represented by an $m\times n$ matrix, and t be the input vector. Without loss of generality, we assume that $A=W^{\dagger}W$ and $f^{\text{forge}}(\textbf{t})$ is defined as:
$f^{\text{forge}}(t_{i})=\begin{cases}0\quad i\leq h\\ t_{i}\quad\text{otherwise},\end{cases}$ (12)
where $t_{i}$ is the $i$-th element of the vector t and $h$ is a positive number.
@@
Since the first $h$ entries of the vector t are replaced with zeros, above condition can be achieved by replacing the corresponding column vectors of the matrix $A$ with zero vectors. Therefore, the entries of $A^{\prime}$ are formulated as
$a^{\prime}_{ij}=\begin{cases}0\quad j\leq h\\ a_{ij}\quad\text{otherwise}.\end{cases}$ (14)
The matrix $A$ is a positive semidefinite matrix, implying that the diagonal entries are non-negative. Moreover, with the entry representation of $A^{\prime}$ in (14), we observe that modifications are only applied to the first $h$ columns, while the rest remain unchanged. Combining the Gershgorin Circle Theorem, we know that the centers of the first $h$ disks of the matrix $A^{\prime}$ are shifted towards zero. Additionally, the radii of all disks, the absolute values of the off-diagonal entries in $A^{\prime}$, are shrunk. Consequently, the upper bound of the largest eigenvalue of the matrix $A^{\prime}$ is tighter compared to that of the original matrix $A$. ∎
It is important to note that the proof presented above is specified for each individual input vector, with each input having its own corresponding matrix $A^{\prime}$ that ensures a bounded Lipschitz constant. This input-specific proof allows for precise theoretical guarantees on robustness. While the corresponding matrix $A^{\prime}$ is auxiliary and used primarily for mathematical proof purposes, there is no need to explicitly compute this matrix during the inference phase.
3.3 Data-driven Algorithm
The choice of a proper $c^{\text{th}}_{i}$ represents a critical balance in our approach. A larger threshold more effectively constrains the adversarial space and reduces the Lipschitz constant. However, excessive restriction can significantly distort the output distribution, potentially degrading overall accuracy on natural examples.
In this paper, we propose a data-driven approach for determining optimal value through the following equation:
@@ 3.3 Data-driven Algorithm
where $\mathcal{S}$ can include all or a subset of images in the training set, $c^{r}$ is a positive number within the range $[0,1]$ and $F_{1\rightarrow i}(x)$ represents the output of the $i$-th layer. Specifically, each layer has its own $c^{\text{th}}_{i}$, but they share the same hyper-parameter $c^{r}$.
Algorithm 1 specifies the implementation details of the forged function. The variable $b$ is used to store the maximum value that appeared in $\mathcal{S}$, as defined in (15), and is initialized during construction. Similar to the implementation of the batchnorm layer, the behavior is depended on the mode configuration. When the mode is set to tracking mode, the variable $b$ is updated accordingly, and the input is set to the output without any modification. Conversely, when the mode is set to inference mode, the value of $b$ is frozen, but the input is updated as defined by (7). By default, the mode is set to inference, and the values of $b$ and $c^{r}$ are zero, respectively. As a result, the mask $x^{\text{mask}}$ are set to $1$, and the algorithm is degraded to the identical function.
Our method offers significant efficiency advantages in practical implementation as the threshold $c^{\text{th}}_{i}$ can be obtained by scanning all data in the set $\mathcal{S}$ once in track mode beforehand. This process requires no gradient computations and adds minimal computational overhead—typically just minutes even on consumer-grade GPUs. Compared to the substantial costs of adversarial training, our approach is essentially cost-free while providing meaningful robustness improvements. Conceptually, the proposed forged function operates analogously to a ReLU function, as it suppresses output values within a specific range, but with a key distinction: the defined range in our forged function is adaptively determined based on the characteristics of the observed dataset rather than being fixed. This data-driven adaptation enables our method to enhance robustness while preserving performance on natural examples.
@@
1:require: Input x, Mode $m$, Hyper-parameter $c^{r}$
@@
5:  for all $x_{i}\in\textbf{x}$ do
6:   if $\text{abs}(x_{i})\leq c^{r}b$ then
7:    $x^{\text{mask}}=f^{\text{masked}}(x)$
8:   else
9:    $x^{\text{mask}}$ = 1
10:   end if
11:   $x_{i}\leftarrow x^{\text{mask}}x_{i}$
12:  end for
13:end if
14:return x